The probability that a player will get 6-11 questions right on a trivia quiz are shown below. (See table)


Table 2: Player Proability
X 6, 7, 8, 9, 10, 11,
P(X) 0.05, 0.1, 0.3, 0.1, 0.15, 0.3
Find the mean. Find the variance. Find the standard deviation.
Answer #1

Anములు, Given Table X 6 7 8 9 10 ll PCXsx) 0.05 0.1 0.3 0.1 0.15 0.3 ? 600.05)+7(0.1) +800-3)+9(0:1) +1000-15) 1. Find the me
Answer #2

SOLUTION :


Mean, m 

= ∑ p(x) * X  

= 0.05*6 + 0.1*7 + 0.3*8 + 0.1*9 + 0.15*10 + 0.3 *11

= 9.1 (ANSWER)


Variance

= ∑ p(x)(m - X)^2 

= 0.05(9.1 - 6)^2 + 0.1(9.1 - 7)^2 + 0.3(9.1 - 8)^2 + 0.1(9.1 - 9)^2 + 0.15(9.1 - 10)^2 + 0.3(9.1 - 11)^2

= 2.49 (ANSWER)


SD 

= sqrt(Variance)

= sqrt(2.49)

= 1.578 (ANSWER).


answered by: Tulsiram Garg